OpenStudy (anonymous):

Math each geometric sequence with the correct sum. a1 = 343, an = -1, r = -1/7 a1 = 3, a8 = 384, r = 2 a1 = 80, n= 7, r = - 1/2 a1 = 5, n= 12, r = 3 a1 = 625, n= 5, r = 3/5 Choose... 215/4 300 1,328,600 630/3 765 1441

OpenStudy (anonymous):

helppppppp

OpenStudy (helder_edwin):

u can use \[ S_n=a_1\cdot\frac{1-r^n}{1-r} \]

OpenStudy (anonymous):

what is the r^n

OpenStudy (helder_edwin):

r is the ratio, for instance, in your first example r=-1/7

OpenStudy (anonymous):



OpenStudy (anonymous):

is that right

OpenStudy (helder_edwin):

in the first you have to find first the number of elements in your sequence

OpenStudy (helder_edwin):

u have a1=343 and r=-1/7 and an=-1

OpenStudy (helder_edwin):

now a2=a1*r a3=a2*r=a1* r^2 then in general an=a1*r^(n-1)

OpenStudy (anonymous):

-1=343*-1/7^(n-1)

OpenStudy (helder_edwin):

yes